OpenStudy (anonymous):

A particle of mass M , initially at rest at x=0, is accelerated by a force that increases in time as F=CT^2. Determine its velocity v as a function of time.

OpenStudy (anonymous):

F(x)=V

OpenStudy (irishboy123):

start with Newton: F = ma. got that? that leaves a = F/m or,..... a(t) = F(t)/m, as you gave a force that varies with time as opposed to the usual Newtonian constant force (and constant acceleration). the same formula holds true BTW, if the Force at time t is F(t), then the acceleration will be F(t)/m. in your case we know that F(t) = ct^2 therefore we know something about the acceleration, don't we? a(t) = ?!?! now, you will need to integrate to switch from acceleration to velocity because d(velocity)dt = acceleration: or S (acceleration) dt = velocity. feel free to let me know how you get on.